2 cars are moving in the same directions with a speed of 30 km/h.They are separated from each other by 5 km.3rd car moving in the opposite direction meets the 2 cars after an interval of 4 minutes.What is the speed of the third car?


Please help me to solve this question.

Grade:Upto college level

5 Answers

Anantha Krishnan
11 years ago

45km/hr.

TANAYRAJ SINGH CHOUHAN
11 years ago

Hi Anurag you posted a question on the two cars moving in same direction with same velocity 30km/hr , but how can the third car meet both cars after the given time interval if both have same velocity and have 5m of initial distance.

Plz. approve my answer by clicking Yes given below, if u loved it... Plz/....

And please reply if the question is to be edited

BEST OF LUCK!

anurag bhattacharjee
11 years ago

Firstly,it,s not my question.It was asked in a test by my teachers.Initially, the cars were moving with a speed of 30 KM/h(in the same direction ) by maintaining a constant difference of 5KM with respect to each other i.e 1st and 2nd car.Secondly, the 3rd car is not meeting the 1st car and 2nd car simultaneously from the opposite direction(since there is sufficient time gap between them).

rupali sehgal
11 years ago

velocity of A wrt C = va-vc

at time t=0 car C met car B

at t=4/60 hrs. car A will meet Car C.

distance traveled by Car A = 5km

=> 5= (Va-Vc)4/60

300/4= 30- Vc

75- 30 = - Vc

Vc = -45 kmph. (- shows opposite direction)

all the best!

 

Vaishali Senad
5 years ago
The two cars are moving in same direction with same speed. So let the speed of two cars be v Km/h and velocity of car moving in opposite direction be u Km/h. So, total speed= total distance/ time= 5/1/15 (4 minutes= 4/60= 1/15 hrs.). So we will get total speed as 75 km/h. Therefore, v+u= 75 I.e., u= 75-30= 45 km/h i.e., the speed of car moving in opposite direction is 45 km/h.